OpenStudy (anonymous):

need help with \[\sqrt{2}(\sqrt{5}-\sqrt{3}-2\sqrt{2})\]

OpenStudy (zzr0ck3r):

\(\sqrt{a}\sqrt{b} = \sqrt{ab}\) distribute

OpenStudy (gorv):

\[\sqrt{2}*\sqrt{5}-\sqrt{3}*\sqrt{2}-2*\sqrt{2}\sqrt{2}=\sqrt{10}-\sqrt{6}-2*\sqrt{4}=\sqrt{10}-\sqrt{6}-4\]

OpenStudy (anonymous):

thank you